OpenStudy (anonymous):

evaluate m/q+np^. for m=24,n=6,p=9 and q=8 114? 489? 580? 2,919?

OpenStudy (aum):

Just plug in the numbers and so the calculation.

OpenStudy (anonymous):

I did and I squared 69 and got 4761 but I don't really know what to go after that like what do I do with the fraction? sorry to bother you also just really horrible at math

OpenStudy (aum):

Something is missing in the problem after np^ (p raised to what?) Assuming it is p^2, we have : m/q + np^2 = 24/8 + 6 * (9)^2 = 3 + 6 * 81 = 3 + 486 = 489. Can you confirm if it is p^2?